OpenStudy (anonymous):

you are filling a pool with water. the pool is a cylinder with a diameter of 12 feet and a height of 8 feet. about how much water will be needed to fill the pool to the very top show all work

OpenStudy (anonymous):

To fill up the whole thing, u need to find the volume: V = 2pi*r*h

OpenStudy (anonymous):

r = d/2 = 12/2 = 6

OpenStudy (anonymous):

r = 6 feet h = 8 feet V = 2pi6*8 = 48*2pi = 96pi

OpenStudy (anonymous):

when bahrom7893 gets here please click on become a fan if this helped, thanks! =)

OpenStudy (bahrom7893):

You will need 96pi ft^3 of water to fill that volume